OpenStudy (mopsy510):

How do you find the dimensions that has something to do with quadratic equations?

OpenStudy (kirbykirby):

Dimensions of what exactly? The question is a little vague. Do you have more context for this?

OpenStudy (mopsy510):

I'll type the whole problem

OpenStudy (mopsy510):

A rectangular painting has dimensions "x" and "x+10". The painting is in a frame 2 in. wide. The total area of the picture and the frame is 900 in. squared. What are the dimensions of the painting?

OpenStudy (mopsy510):

Hello?

OpenStudy (kirbykirby):

 So since along each side, you have (2+2)=4 inches LESS for the painting, each side will have dimensions of the painting would be "x-4" and "x+6" (since you are doing x+10-4) So the area = side*side = (x-4)(x+6) We know the are is 900 so, 900 = (x-4)(x+6) = \(x^2+6x-4x-24=x^2+2x-24\) so, \(0=x^2+2x-924\) ... and use the quadratic equation to find x